# OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion

### v1

@@ Abstract
There has been significant progress in open-source text-only translation large language models (LLMs) with better language coverage and quality. However, these models can be only used in cascaded pipelines for speech translation (ST), performing automatic speech recognition first followed by translation. This introduces additional latency, which is particularly critical in simultaneous ST (SimulST), and prevents the model from exploiting multimodal context, such as images, which can aid disambiguation. Pretrained multimodal foundation models (MMFMs) already possess strong perception and reasoning capabilities across multiple modalities, but generally lack the multilingual coverage and specialized translation performance of dedicated translation LLMs. To build an effective multimodal translation system, we propose an end-to-end approach that fuses MMFMs with translation LLMs. We introduce a novel fusion strategy that connects hidden states from multiple layers of a pretrained MMFM to a translation LLM, enabling joint end-to-end training. The resulting model, OmniFusion, built on Omni 2.5-7B as the MMFM and SeedX PPO-7B as the translation LLM, can perform speech-to-text, speech-and-image-to-text, and text-and-image-to-text translation. Experiments demonstrate that OmniFusion effectively leverages both audio and visual inputs, achieves a 1-second latency reduction in SimulST compared to cascaded pipelines and also improves the overall translation quality [^1] Code is available at https://github.com/saikoneru/OmniFusion.
OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion
@@ 1 Introduction
In recent years, research in natural language processing (NLP) has increasingly shifted from unimodal foundation models (Touvron et al., 2023; Grattafiori et al., 2024) to multimodal foundation models (MMFMs) (Abdin et al., 2024; Goel et al., 2025; Xu et al., 2025), as many tasks benefit from contextual information derived from multiple sources. Although unimodal systems can perform strongly in certain settings, further progress also requires models to be able to take advantage of additional contextual cues without degrading quality.
One representative task that can benefit from such multimodal integration is speech translation (ST), where the goal is to translate spoken input in a source language into a target language. Over the years, advances in data scale, modeling techniques, and architectural design have led to substantial improvements in ST quality across many language pairs and domains (Ahmad et al., 2024; Agostinelli et al., 2025). However, the use of multimodal context—such as visual information from images or presentation slides accompanying speech—remains relatively underexplored (Liu and Niehues, 2024; Gaido et al., 2024a; Sinhamahapatra and Niehues, 2025).
@@
Although MMFMs are capable of processing multiple modalities jointly, their multilingual translation capabilities are generally weaker than those of specialized translation large language models (LLMs) (Alves et al., 2024; Cheng et al., 2025). As a result, directly using MMFMs for ST often yields inferior translation quality compared to approaches that rely on strong text-based translation components (Appendix: Table 4). They also do not cover several languages compared to specialized translation LLMs. Consequently, a practical solution is to adopt a cascaded pipeline, where automatic speech recognition (ASR) is first performed, followed by machine translation (MT). This modular setup allows each component to be optimized independently and integrated during inference (Ahmad et al., 2024; Agostinelli et al., 2025; Koneru et al., 2025). However, since the translation model operates solely on text, it cannot exploit contextual information from the original speech or visual input. Moreover, the cascaded design introduces additional generation steps, leading to increased latency compared to end-to-end (E2E) models—an issue that is particularly critical for simultaneous ST (SimulST) applications (Agostinelli et al., 2025; Papi et al., 2025a).
Therefore, we aim to develop an E2E MMFM that excels in multimodality, multilinguality and efficient for SimulST. This can be achieved by extending the capabilities of a translation LLM to handle additional modalities. Few prior works have explored adding audio or image inputs to multilingual LLMs (Ambilduke et al., 2025; Viveiros et al., 2025). However, these approaches typically rely on training from scratch or only using pretrained audio/vision encoders and do not leverage the higher-level representational capabilities of existing MMFMs. Furthermore, training is ineffective as the base LLM needs to learn the modality mapping from scratch.
In this work, we propose leveraging pretrained MMFMs to introduce multimodal capabilities into a multilingual LLM with a E2E architecture. Rather than transferring only vision or audio encoders, we extract and utilize hidden-state representations from multiple layers of the MMFM to capture both the perceptual and reasoning components of multimodal understanding. Specifically, we adapt the translation LLM by incorporating hidden states from the first, middle, and last layers of the MMFM, thereby enabling the transfer of both low-level perception and high-level semantic understanding.
To further strengthen cross-modal alignment, we introduce Optical Character Recognition (OCR) as a bridge task for the image modality, analogous to how ASR serves as an alignment mechanism for audio. This strategy allows the model to effectively leverage pretrained multimodal representations while receiving direct cross-modal supervision.
We use Qwen Omni 2.5 7B (Xu et al., 2025) as the MMFM and SeedX 7B (Cheng et al., 2025) as the translation LLM to validate our approach on SimulST and multimodal MT tasks. We refer to the resulting fused model as OmniFusion.
Our main contributions are summarized as follows:
  • We propose a gated fusion approach that extracts multimodal embeddings from MMFMs, allowing translation LLMs to fully utilize the pretrained perceptual and interpretive capabilities of MMFMs.
  • We show the benefit of using slides for context and achieve an improvement in latency of approximately $1$ second in the SimulST scenario compared to cascaded fine-tuned models while having better quality. (Figure 2).
  • We observe a reduction in the major and critical errors for offline ST through self-cascading and effective use of the image context (Table 2).
  • We achieve state-of-the-art performance in the CoMMuTE benchmark for image-text translation highlighting the ability of the model to exploit images (Table 3).
@@ 2 Approach
We propose to fuse a MMFM with a task-specific translation LLM to leverage their complementary strengths and enable E2E training for improved latency. Furthermore, we aim to incorporate not only the perception capabilities of the MMFM but also its multimodal reasoning and understanding.
First, we describe the overall architecture and training process of our approach, considering a general input and output sequence independent of the specific task. We formally define the information flow, training loss, and the fusion of MMFM layers. Next, we detail the training tasks used to align multimodal representations from both image and audio modalities to the translation LLM. Finally, we discuss the training prompts that enable a self-cascading inference mode, allowing the model to enhance translation quality at the cost of increased latency. An overview of the architecture and supported tasks is in Figure 1.
@@ 2.1 Architecture
Let us denote the specialized translation LLM as $\mathcal{M}_{\text{NMT}}$ and the MMFM as $\mathcal{M}_{\text{MMFM}}$, where both use a language model as their backbone. The $\mathcal{M}_{\text{NMT}}$ is trained for translation via supervised fine-tuning with the following loss function:
@@
For $\mathcal{M}_{\text{MMFM}}$, the image and audio modalities are first processed by modality-specific encoders and then projected to match the embedding dimension of $\mathcal{M}_{\text{MMFM}}$, denoted as $D_{\text{MMFM}}$. This projection allows the encoder outputs to be concatenated as a sequence of tokens, enabling the MMFM to operate similarly to a standard LLM, such as $\mathcal{M}_{\text{NMT}}$.
To fuse both the perception and higher-level understanding of the MMFM, we hypothesize that lower layers predominantly capture perceptual features, while higher layers encode more abstract reasoning as computation progresses. Accordingly, we extract the first, middle, and last hidden layers from the MMFM, denoted as:
$h^{(1)}_{t}$ $=\text{MMFM}_{\text{layer 1}}(x_{t}),$ (2)
$h^{(\text{mid})}_{t}$ $=\text{MMFM}_{\text{layer mid}}(x_{t}),$
$h^{(\text{last})}_{t}$ $=\text{MMFM}_{\text{layer last}}(x_{t}),$
$h^{(1)}_{t},\,h^{(\text{mid})}_{t}$, $\,h^{(\text{last})}_{t}\in\mathbb{R}^{L_{m}\times D_{\text{MMFM}}},t=1,\dots,L_{m}$
where $L_{m}$ is the length of the MMFM hidden states.
Directly concatenating these multi-dimensional representations into $\mathcal{M}_{\text{NMT}}$ poses three challenges. First, the sequence length triples, significantly increasing computational and memory requirements. Second, the relevance of each layer may vary depending on the task, necessitating selective weighting. Third, the embedding dimension of $\mathcal{M}_{\text{MMFM}}$ may not match that of $\mathcal{M}_{\text{NMT}}$.
To overcome these challenges, we introduce a gated fusion layer that determines the contribution of each hidden state along the sequence length. Each hidden state is multiplied by its corresponding gate values, summed across layers, and then projected through an MLP to align the embedding dimension with that of the translation model:
@@
where $D_{\text{Trans}}$ is the embedding dimension of $\mathcal{M}_{\text{NMT}}$ and $m_{t}$ is the fused multimodal representation at time step $t$.
The fused multimodal embeddings $m_{t}$ can now be treated like additional token embeddings and concatenated with the $\mathcal{M}_{NMT}$ token embeddings:
@@ 2.2 Training
We have shown how MMFM internal representations can be extracted and integrated into our downstream model. However, leveraging these representations effectively also requires a training process that explicitly aligns multimodal features so that the LLM can exploit them. The exact alignment procedure depends on the downstream tasks we expect the fused model to perform after fine-tuning.
In this work, our goal is to obtain a translation model that can utilize both audio and image context. Therefore, we must make the translation LLM context-aware by training to exploit these representations. We consider $3$ translation tasks in this work and are defined as follows:
  • Speech Translation (ST): Generate a translation in the target language given only audio in the source language.
  • Speech-Image Translation (SIT): Generate a translation in the target language given source-language audio together with an image that provides additional context (e.g., a slide corresponding to the spoken segment).
  • Text-Image Translation (TIT): Generate a translation in the target language given the source text and an aligned image, such as in caption translation scenarios.
While we can train the model to perform translation directly, decomposing the task into intermediate steps can be beneficial. For example, translating English audio to German can be approached by first predicting ASR outputs and then producing the translation by continuing generation conditioned on the ASR hypothesis. This decomposition not only simplifies learning but also provides explicit alignment signals that help map audio representations into the LLM’s semantic space.
When vision input is available—particularly for slide-based scenarios—we additionally introduce OCR as a bridge task before generating the final translation. In this case, the model must predict both OCR text and the final translation using the same audio–visual representation. This encourages alignment between the modalities and explicitly informs the model that the visual context is meaningful and should be utilized.
During training, we stochastically decide—based on a specified probability for each data source—whether an example should be trained in self-cascading mode and along which modality the cascade should occur. Furthermore, we provide modality-specific prompt to the MMFM for handling different modality combination scenarios. Pseudocode for building the prompt for each task during training is shown in Algorithm˜1.
We also freeze the MMFM to preserve its pretrained abilities while fine-tuning the translation LLM with LoRA (Hu et al., 2021). We use LoRA for computational benefits and not go Out-Of-Memory during fine-tuning.
@@ 3.1 Training Data
We aim to train the fused model to perform translation tasks across both speech and image modalities. An overview of the data sources is provided in Table˜1.
For ST, we use subsets of Europarl-ST (Iranzo-Sánchez et al., 2020) and Covost-2 (Wang et al., 2020), sampling across multiple target languages. We limit training data to these subsets due to computational constraints and to demonstrate that full datasets are not strictly necessary.
To enable the model to exploit images when translating audio, we use the M3AV corpus (Chen et al., 2024), which contains English scientific talks paired with corresponding slides. The slides also include Paddle-OCR (Cui et al., 2025) predictions, allowing the model to be trained either to predict translations directly or via an intermediate OCR step for better alignment.
Since M3AV has limited language coverage, we augment the data by synthetically translating ASR transcripts into additional target languages (M3AV-Aug). This enables the model to handle ST with slides and supports the self-cascade mode, performing ASR first and then translation.
Finally, for TIT, we employ the Multi30k dataset (Elliott et al., 2016), which contains images paired with captions in multiple languages. This task encourages the model to go beyond OCR and to leverage visual content, including objects and scene understanding, for translation.
@@ 3.2 Models
Several MMFMs have been developed rapidly in recent years, showing significant improvements in both quality and efficiency (Abdin et al., 2024; Team et al., 2025). In our scenario, we require a model capable of processing both audio and image modalities, while also supporting cross-modal abilities. Accordingly, we select Qwen Omni 2.5B [^2] Qwen/Qwen2 Audio.5-Omni-7B (Xu et al., 2025), due to its strong multimodal capabilities and the potential to incorporate video in future experiments.
For the translation LLM, we choose the recently proposed SeedX PPO 7B [^3] ByteDance-Seed/Seed-X-PPO-7B (Cheng et al., 2025), which provides broad language coverage and strong translation performance compared to closed-source alternatives.
We use the training data mentioned above and denote the fine-tuned model as OmniFusion for the rest of the paper. While we report results using these specific models, it is important to note that our approach is not limited to these base LLMs. The proposed fusion strategy does not require massive pre-training or fine-tuning resources and can be adapted in more modest computing environments, such as university clusters. The training hyper-parameters are specified in Table˜10. For validation, we used Covost-2 validation data and select the checkpoint with the least loss in 20k steps.
Dataset # Examples Language Modality Task
M3AV 122k en Speech, Image ASR, OCR
M3AV - Aug 122k
zh, fr, de, it,
ja, ko, pt, ru,
es, vi, ar, cs,
hr, da, nl, fi,
hu, id, ms, nb,
no, pl, ro, tr
Speech, Image ST, OCR
Europarl 260k*
de, es, fr, it, nl,
pl, pt, ro
Speech ST
Covost2 de, ja, tr, zh Speech ST
Coco-Captions 87k cs, de, fr Image TIT
Table 1: Overview of data sources collected for E2E training. * We sample together 260k from the concatenated data uniformly across the language pairs. All the speech and text data are only available with English as the source language. Language code mapping is presented in Table˜10.
@@ 3.3 Evaluation: Data & Metrics
For evaluating the performance of multimodal Simul-ST systems, we require aligned audio-visual translation data. Additionally, we avoid using M3AV for testing, as it overlaps with our training data, monolingual and we aim to measure the model’s robustness on unseen domains.
Consequently, we use the MCIF test set (Papi et al., 2025b), which contains video and audio recordings of ACL talks in three target languages from English. We report latency with Average Lagging (Ma et al., 2020) and translation quality results on this dataset.
For TIT, we focus on test cases where the image is crucial for disambiguating the source sentence. For this, we use the CoMMuTE dataset (Futeral et al., 2023, 2025), which provides a small but multilingual targeted TIT test set.
We measure translation quality using XCOMET-XL (Guerreiro et al., 2024) for ST scenarios, as it offers high-quality evaluation and allows analysis of minor, major, and critical errors. For TIT, we use the COMET [^4] Unbabel/wmt22-comet-da model (Rei et al., 2022) directly, without providing error spans, to evaluate aligned references.

### v2

# OmniFusion: Simultaneous Multilingual and Multimodal Translations via Modular Fusion

@@ Abstract
Open-source text-only translation Large Language Models (LLMs) have rapidly improved in multilingual coverage and translation quality, but their unimodal design limits their applicability to multimodal translation. In speech translation (ST), they typically operate in cascaded pipelines with automatic speech recognition followed by translation, introducing additional latency particularly for simultaneous ST (SimulST) and they cannot exploit visual context for disambiguation in either speech or caption translation. In contrast, pretrained multimodal foundation models (MMFMs) provide strong cross-modal perception and reasoning but lack the multilingual depth and translation performance of specialized translation LLMs. We introduce OmniFusion, a unified architecture that combines a multimodal foundation model with a translation-specialized LLM through a lightweight gated fusion module. Built with Qwen2.5-Omni-7B as the MMFM and Seed-X-PPO-7B as the translation LLM, OmniFusion supports speech-to-text, speech-and-image-to-text, and text-and-image-to-text translation within a single architecture. Experiments show that OmniFusion effectively leverages audio and visual cues and reduces SimulST latency by one second compared to cascaded pipelines while maintaining competitive translation quality. [^1] Code is attached as a zip file.
OmniFusion: Simultaneous Multilingual and Multimodal Translations via Modular Fusion
@@ 1 Introduction
In recent years, NLP research has shifted from unimodal foundation models (Touvron et al., 2023; Grattafiori et al., 2024) to multimodal foundation models (MMFMs) (Abdin et al., 2024; Goel et al., 2025; Xu et al., 2025), as many tasks benefit from contextual information across sources. While unimodal systems can perform well in certain settings, further progress requires models that exploit additional cues without degrading quality.
Translation is a representative task where multimodal integration is beneficial. In Speech Translation (ST), the goal is to convert spoken input in a source language into a target language (Ahmad et al., 2024; Agostinelli et al., 2025). Despite recent advances, leveraging visual context such as images or presentation slides accompanying speech remains underexplored (Liu and Niehues, 2024; Gaido et al., 2024a; Sinhamahapatra and Niehues, 2025). Similarly, caption translation involves short, image-linked text, where visual information can aid disambiguation and grounding.
@@
Although MMFMs can process multimodal inputs, their multilingual translation performance is generally weaker than that of specialized translation LLMs (Alves et al., 2024; Cheng et al., 2025), leading to lower translation quality and limited language coverage. In ST, a common approach is a cascaded pipeline: automatic speech recognition (ASR) followed by machine translation (MT) (Ahmad et al., 2024; Agostinelli et al., 2025; Koneru et al., 2025). While this modular approach allows independent optimization, it prevents the translation model from leveraging multimodal context and adds latency, a limitation especially critical in simultaneous ST (SimulST) scenarios (Agostinelli et al., 2025; Papi et al., 2025a).
These limitations motivate the development of an end-to-end multilingual and multimodal translation system supporting ST, image-aware ST, and caption translation. A unified model can exploit shared data across tasks and simplify deployment. Extending a translation LLM to handle additional modalities is one approach, but prior work (Ambilduke et al., 2025; Viveiros et al., 2025) often trains from scratch or relies solely on pretrained encoders, without fully leveraging the higher-level representations of existing MMFMs, making training inefficient.
We propose leveraging pretrained MMFMs to introduce multimodal capabilities into a multilingual LLM within a unified E2E architecture. Our approach tightly couples the two models during training and employs a multi-task objective augmented with auxiliary tasks, including OCR and ASR with self-cascading, to enhance multimodal alignment and translation performance.
Our contributions are threefold: (1) We present OmniFusion, an E2E model built using translation LLMs and pretrained MMFMs that jointly supports speech, text and image translation tasks. (2) We introduce a gated fusion architecture that selectively integrates representations from multiple MMFM layers into a translation LLM, enabling efficient training and inference for context-aware translations; and (3) We demonstrate that OmniFusion reduces SimulST latency by $\sim 1$ second (Figure 3), lowers major and critical error rates for offline ST (Table 1), and achieves state-of-the-art results on CoMMuTE (Table 2).
OmniFusion demonstrates that MMFMs and translation LLMs can be effectively combined within a single architecture for multilingual multimodal translation.
@@ 2 Approach
We propose fusing a MMFM with a task-specific translation LLM to leverage their complementary strengths. The MMFM provides perceptual representations, while the translation LLM provides multilingual understanding and generation capabilities [^2] Using the MMFM as encoder does not provide significant computational overhead as decoders are primary latency bottleneck (Guta et al., 2025)..
We first describe the overall architecture and training process, abstracting over task-specific inputs and outputs. Motivated by a probing experiment, we define the information flow, training loss, and fusion of MMFM layers. We then detail training tasks aligning image and audio modalities to the translation LLM and discuss prompts enabling self-cascading inference, which improves translation quality at higher latency. An overview is shown in Figure 1.
@@ 2.1 Architecture
Let $\mathcal{M}_{\text{NMT}}$ be the translation LLM and $\mathcal{M}_{\text{MMFM}}$ the MMFM, both with a language model backbone. $\mathcal{M}_{\text{NMT}}$ is trained via supervised fine-tuning with:
@@
For $\mathcal{M}_{\text{MMFM}}$, image and audio inputs are encoded and projected to the MMFM embedding dimension $D_{\text{MMFM}}$, forming a token sequence compatible with the LLM.
2.1.1 Probing MMFM Representations
Figure 2: Average cosine similarity between hidden representations of adjacent layers, computed with Qwen-Omni 2.5 7B on the MCIF test set (en $\rightarrow$ de) under different modality conditions.
To capture both perceptual and higher-level understanding, we hypothesize that the lower MMFM layers encode perceptual features and higher-level abstract reasoning. We validate this by computing cosine similarity between consecutive layer outputs, which is commonly used as a proxy for layer importance in model pruning literature. We compute with different modalities and visualize in Figure˜2. We see that the first and last layers exhibit the largest changes, while other layers provide small updates. Based on this, we propose using the first, middle, and last layers as inputs to generate multimodal representations as they capture the three regions [^3] We perform ablation using all layers in Table 3 and empirically validate this choice..
2.2 Gated Fusion
Accordingly, we extract these layers from the MMFM, denoted as:
$h^{(k)}_{t}$ $=\text{MMFM}_{\text{layer }k}(x_{t}),\quad k\in\{1,\text{mid},\text{last}\},$ (2)
$h^{(k)}_{t}$ $\in\mathbb{R}^{L_{m}\times D_{\text{MMFM}}},\quad t=1,\dots,L_{m}$
where $L_{m}$ is the length of the MMFM hidden states. Directly concatenating these multi-dimensional MMFM representations into $\mathcal{M}_{\text{NMT}}$ is problematic because it triples the sequence length, layer relevance may vary by task, and embedding dimensions may not match. To address this, we use a gated fusion layer that weights each hidden state, sums across layers, and projects the result via an MLP to align with the translation model:
@@ 2.2 Gated Fusion
where $\oplus$ is concatenation done per each token across layers, $D_{\text{Trans}}$ is the embedding dimension of $\mathcal{M}_{\text{NMT}}$ and $m_{t}$ is the fused multimodal representation at time step $t$. The fused multimodal embeddings $m_{t}$ can now be treated like additional token embeddings and concatenated with the $\mathcal{M}_{NMT}$ token embeddings:
@@ 2.3 Training
Effectively leveraging MMFM representations requires aligning multimodal features with the translation LLM’s token representations. We achieve this through multitask training that includes ST, ASR, and OCR objectives.
2.3.1 Multi-task Learning
We train the model on three primary multimodal translation tasks. In Speech Translation (ST), the model generates a target-language translation from source-language audio. In Speech–Image Translation (SIT), it translates audio together with an image providing additional context, such as a presentation slide aligned with the spoken segment. In Text–Image Translation (TIT), the model translates source text paired with an aligned image, as in caption translation scenarios. Training across these tasks exposes the model to different modality combinations, enabling effective alignment and exploitation of multimodal representations.
2.3.2 Self-Cascading (ASR and OCR)
In addition to multi-task learning, we adopt self-cascading, which decomposes translation into intermediate steps. For English-to-German ST, the model first generates the ASR transcription, then produces the final translation conditioned on it, providing explicit alignment between audio representations and the LLM’s semantic space. For visual inputs, OCR text serves as an intermediate step before translation, aligning visual and textual representations. During training, self-cascading is applied stochastically based on task and modality, with modality-specific prompts (Algorithm˜1). We freeze the MMFM and fine-tune the translation LLM with LoRA (Hu et al., 2021), reducing compute costs while retaining pretrained multimodal capabilities. We also find that adding LoRA to the MMFM degrades performance due to catastrophic forgetting.
@@ 3.1 Training Data
We train the fused model to perform translation across speech and image modalities. An overview of data sources is provided in Appendix Table˜6. For ST, we use subsets of Europarl-ST (Iranzo-Sánchez et al., 2020) and Covost-2 (Wang et al., 2020), sampling multiple target languages. Training is limited to these subsets to reduce computational costs. To exploit images in audio translation, we use the M3AV corpus (Chen et al., 2024), containing English scientific talks paired with slides. The slides include Paddle-OCR (Cui et al., 2025) predictions, allowing training either for direct translation or via an intermediate OCR step for better alignment. Since M3AV has limited language coverage, we augment ASR transcripts with synthetic translations (M3AV-Aug), enabling ST with slides and supporting self-cascade training.
For TIT, we use Multi30k (Elliott et al., 2016), which pairs images with multilingual captions. This task encourages leveraging visual content beyond OCR, including objects and scene understanding, to improve translation.
@@ 3.2 Models
Several MMFMs have recently shown strong gains in quality and efficiency (Abdin et al., 2024; Team et al., 2025). As our setting requires processing both audio and image modalities with cross-modal reasoning, we use Qwen Omni 2.5 7B [^4] Qwen/Qwen2.5-Omni-7B (Xu et al., 2025). For translation, we adopt SeedX PPO 7B [^5] ByteDance-Seed/Seed-X-PPO-7B (Cheng et al., 2025), which provides broad language coverage and strong translation performance compared to closed-source alternatives. Using the training data described above, we refer to the resulting fine-tuned model as OmniFusion. While we report results with these models, the approach is general and applicable to other base LLMs. Training hyperparameters are listed in Table˜14. Validation uses Covost-2 dev data, selecting the checkpoint with the lowest loss after 20k steps.
@@ 3.3 Evaluation: Data & Metrics
Evaluating multimodal SimulST requires aligned audio–visual translation data. We exclude M3AV from testing due to training overlap and its monolingual nature, and instead assess robustness on unseen domains using the MCIF test set (Papi et al., 2025b), which contains ACL talk recordings translated from English into three target languages, and FLEURS (Conneau et al., 2023) for multilingual coverage. Latency is reported using Average Lagging (Ma et al., 2020), and translation quality is measured on these datasets.
For TIT, we target cases where visual context is crucial for disambiguation, using the CoMMuTE benchmark (Futeral et al., 2023, 2025). Translation quality is evaluated with XCOMET-XL (Guerreiro et al., 2024) for ST, enabling detailed error analysis, and with COMET [^6] Unbabel/wmt22-comet-da (Rei et al., 2022) for TIT.